OpenStudy (compassionate):

f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)) So, lets see... -4x(10x - 6) + 7, right?

OpenStudy (shamil98):

No, you replace the x in -4x + 7 with the (10x-6).

OpenStudy (compassionate):

What is all this business about replacing x. WHAT X?

OpenStudy (shamil98):

f(x) = -4x + 7 you see the x right next to the -4, THAT X.

zepdrix (zepdrix):

\[\Large\rm f(\color{royalblue}{x})=-4\color{royalblue}{x}+7\]\[\Large\rm f(\color{royalblue}{g(x)})=-4\color{royalblue}{g(x)}+7\]\[\Large\rm f(\color{royalblue}{g(x)})=-4\color{royalblue}{(10x-6)}+7\]

zepdrix (zepdrix):

Yah that x shouldn't next to the -4 anymore, after you've replaced it :o

OpenStudy (compassionate):

Oh, now I see, Zepdrix. So f(g(x)) = -4(10x - 6) + 7 -40x +24 + 7 Which yields.... -40x + 31 What is the next step, your heinous?

zepdrix (zepdrix):

That would be as far as you can simplify it :) See the notation on the left side of the equals? f(g(x)), it's still a function of x at this point, it's not being evaluated at any particular number.

zepdrix (zepdrix):

For example if you wanted to evaluate this composition at x=2, you would have, f(g(2)) = -40(2)+31 f(g(2)) = -49 But that's not what they asked for here :)

OpenStudy (compassionate):

What is this "composition" you speak of? And how do I know what I need to "evaluate."

zepdrix (zepdrix):

So we started with f(x) and g(x). \(\Large\rm f(g(x))\) reads -> f of g of x sometimes we use this notation instead (meaning the same thing): \(\Large\rm (f \circ g)(x)\) reads -> f composed of g of x It's a composition of functions, a function within a function.

zepdrix (zepdrix):

The question said \(\Large\rm find~f(g(\color{royalblue}{x}))\) Pay attention to the blue x, if it's a number, then that's when they're asking you to evaluate it at a particular number.

zepdrix (zepdrix):

If it's a number, you should end up with a number after doing your calculations. If it's a variable (like they asked for), you should end up with a function (stuff including x, like we did).

OpenStudy (compassionate):

Okay, so since we have "f(g(x))" We're given "x" which is what we're being asked to find. For example, if we had f(g(1)), we'd look for 1, right?

OpenStudy (compassionate):

So now that I have: -40x + 31 I plug that into f(x) f(-40x + 31) = -4(-40x + 31) + 7 Correct?

zepdrix (zepdrix):

If we were asked to find f(g(1)), we could do the same steps we did, first finding f(g(x)), f(g(x)) = -40x + 31 and then plug x=1 into the function to find f(g(1)) f(g(1)) = -40(1) + 31

OpenStudy (compassionate):

But we need to plug -40x + 31 into f(x), not 1, right?

zepdrix (zepdrix):

What you did was, you plugged f(g(x)) into your original f(x) function. So what you just found was f(f(g(x))). Lemme do it in color, maybe it'll be a little more clear.

zepdrix (zepdrix):

We found \(\Large\rm f(g(x))\), and you plugged that whole thing into \(\Large\rm f\left(\color{orangered}{x}\right)=-4(\color{orangered}{x})+7\) and it gave you \(\Large\rm f\left(\color{orangered}{f(g(x)}\right)=-4(\color{orangered}{f(g(x)})+7\) \(\Large\rm f\left(\color{orangered}{f(g(x)}\right)=-4(\color{orangered}{-40x+31})+7\)

zepdrix (zepdrix):

That's a little too fancy :O That's not what we want to be doing hehe

OpenStudy (compassionate):

Well, what are we trying to do cx

zepdrix (zepdrix):

So you're feeling ok with this first step, yes? Plugging a function into a function? At least a little bit better?